OpenStudy (anonymous):

Jane poured 12 cups of water into a rectangular container measuring 12 inches by 8 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (anonymous):

@tester97 i need you one more time

OpenStudy (anonymous):

i think wolf trollin he been replying for the past 20 mins

OpenStudy (wolf1728):

No I'm typing - - please let me finish

OpenStudy (wolf1728):

A) 12inches *8inches *9centimeters = 12*2.54*8*2.54*9 = 5,574.18 cc B) 12 cups = 12*14.44 = 173.28 cubic inches 2.54 * 2.54* 2.54 = 16.387064 cubic centimeters per cubic inch There are 16.387064 cc per cubic inch so 12 cups = 173.28 * 16.38704 = 2,839.55 cc were added Since you are so impatient, I'll let yu or someone else determine C. :-)

OpenStudy (anonymous):

sorry i just needed this last question

OpenStudy (anonymous):

and i was stuck on it

OpenStudy (wolf1728):

well is it right or an I wrong?

OpenStudy (anonymous):

ill have to see

OpenStudy (anonymous):

i gotta do c now then we will know

OpenStudy (wolf1728):

Okay well you gave me a medal so I'll try C if you want

OpenStudy (anonymous):

go for it cause i was just gunna turn in a and ab and forget about c casue im stumped

OpenStudy (wolf1728):

9 cm must be the lowest point. whole container = 5,574.18 cc 2,839.55 cc were added SO, container was 2,839.55 /5,574.18 0.5094112497 full Container filled to 9 cm * 0.5094112497 = 4.5847012475 centimeters

OpenStudy (anonymous):

thnks

OpenStudy (wolf1728):

Wow - that had a LOT of tricks to it - I was messing it up all over the place